# Language Model Preference Evaluation with Multiple Weak Evaluators

### v3

@@ Appendix I Expanded Related Work on Preference Denoising for LLMs
The selection of weak evaluators is a critical factor influencing the performance of GED. While our initial experiments demonstrated the benefits of combining weak evaluators, further analysis is necessary to understand how their diversity and capabilities affect GED’s effectiveness. To address this, we conducted additional experiments evaluating three different configurations of evaluator sets: the original set comprising Llama3-8B, Mistral-7B, and Qwen2-7B; a diverse set that adds Gemma-9B [^5] https://huggingface.co/google/gemma-2-9b-it to introduce more variation in model architecture and training data; and a higher-capability set that replaces smaller models with larger ones, specifically Llama3-70B, Mistral-8×7B, and Qwen2-72B. The results, shown in Table 9, reveal two key insights. First, incorporating diversity by adding Gemma-9B improves GED’s performance slightly, increasing the average score from 56.05% to 56.39%. This suggests that models with diverse training data and architectures can contribute to more robust aggregated evaluations. Second, replacing smaller evaluators with higher-capability models yields a more substantial improvement, with the average score rising to 60.89%. Notably, benchmarks like AlpacaEval and GAIA benefit the most from the advanced reasoning and language understanding capabilities of larger models. These findings demonstrate the importance of both diversity and capability in evaluator selection. Diversity provides marginal gains by bringing varied perspectives, while higher-capacity models contribute to more significant improvements by enhancing the overall quality of evaluations. This suggests that, when computational resources allow, incorporating advanced models into the evaluator set can meaningfully boost GED’s performance.
@@
Table 9: Performance comparison with different evaluator sets across five benchmarks. The results highlight the impact of evaluator diversity and capability on GED’s effectiveness.
@@
Appendix K Cost Considerations and Active Learning with ActiveGED
Using multiple evaluators and aggregating their preferences can be computationally expensive, especially when constructing dense preference graphs. This limitation becomes more pronounced in scenarios where preferences across all pairs of evaluators are required, which scales quadratically with the number of responses or models being compared. To address this, we clarify the computational trade-offs and propose an active learning-based approach to reduce the number of required pairwise evaluations. To further reduce the number of pairwise evaluations needed, we developed an active learning algorithm called ActiveGED. This algorithm strategically selects the most informative pairs for evaluation, effectively lowering the overall computational cost while maintaining high performance. ActiveGED combines random sampling with uncertainty-based selection to maximize information gain from each additional pairwise evaluation. We evaluated ActiveGED under budget constraints of 30% and 50% of the total possible pairwise comparisons, where we set the random budget ratio as 0.5. The results, presented in Table 10, demonstrate that ActiveGED achieves competitive performance with significantly fewer evaluations. For example, under a 50% budget, ActiveGED retains most of the performance benefits of full GED while cutting the number of pairwise comparisons in half.
@@
Evaluators Set of GED HumanEval AlpacaEval MATH GSM8k GAIA Avg.
@@
ActiveGED demonstrates that by carefully selecting the most informative pairs, it is possible to achieve competitive performance with significantly fewer evaluations. This makes GED more scalable and practical for real-world applications where computational resources are limited.
Appendix L Mitigating Evaluator Biases in GED
Evaluator biases, such as position bias and token bias, can significantly impact the accuracy and fairness of the preference graphs used in GED. Addressing these biases is crucial for ensuring reliable evaluations and robust performance. In this section, we describe the strategies employed in our framework to mitigate these biases and discuss potential areas for future improvement. Position bias arises when evaluators exhibit a preference for a particular position in a pairwise comparison, such as consistently favoring the first or second option regardless of content. To counter this, we explicitly include both orderings of each question and its candidate answers in the response selection setting. Specifically, for a question $Q$ with two candidate answers $A_{1}$ and $A_{2}$, we evaluate both configurations:
@@
By testing both (Q, A1, A2) and (Q, A2, A1), we ensure that any positional preferences of the evaluators are balanced out when constructing the preference graph. This approach minimizes the impact of position bias on the overall rankings. Token bias occurs when an evaluator exhibits a latent preference for a specific option or label (e.g., consistently favoring ”Option A” over ”Option B”). Our framework implicitly mitigates token bias through the aforementioned position-swapping strategy, which prevents evaluators from associating a fixed label with a specific position. By averaging the results across both orderings, any systematic preference for a particular label is effectively neutralized.
Appendix M Denoising of Preference Graphs for GED
In this section, we describe the denoising procedure used in the GED framework, specifically for transforming the aggregated preference graph into a DAG. Let $G=(V,A)$ be a simple connected directed graph, where $n=|V|$ and $m=|A|$, with $n-1\leq m\leq\binom{n}{2}$. A feedback arc set (FAS) of $G$, denoted as $R(G)$, is a set of arcs whose removal transforms $G$ into a DAG. The minimum feedback arc set $R^{*}(G)$ is the FAS of minimum cardinality, and finding $R^{*}(G)$ is the well-known FAS problem.
Consider a scenario where the vertices of $G$ are arranged sequentially along a horizontal line and labeled as $v_{1},v_{2},\ldots,v_{n}$ from left to right. This arrangement is referred to as a vertex sequence and is denoted by $s=v_{1}v_{2}\ldots v_{n}$. Each vertex sequence $s$ induces a feedback arc set $R(s)$, consisting of all leftward arcs $v_{j}\rightarrow v_{i}$ for $j>i$. The FAS problem can therefore be reformulated as finding a vertex sequence $s^{*}$ such that $R(s^{*})=R^{*}(G)$. Our proposed algorithm, GED, computes a vertex sequence $s$ that corresponds to a minimal feedback arc set $R(s)$. The algorithm iteratively removes vertices (and their incident arcs) from $G$, focusing on sinks, sources, and vertices that maximize a specific property. For any vertex $u\in V$, let $d(u)$ denote its degree, $d^{+}(u)$ its outdegree, and $d^{-}(u)$ its indegree, such that $d(u)=d^{+}(u)+d^{-}(u)$. At each step, after removing sinks and sources, the algorithm selects a vertex $u$ for which $\delta(u)=d^{+}(u)-d^{-}(u)$ is maximized. If the removed vertex $u$ is a sink, it is concatenated with a vertex sequence $s_{2}$; otherwise, it is concatenated with $s_{1}$. Once $G$ is reduced to an empty graph, the final vertex sequence $s$ is obtained by concatenating $s_{1}$ and $s_{2}$. The detailed steps of the algorithm are shown in Algorithm 4. In our GED framework, this denoising step is essential for ensuring that the aggregated preference graph becomes acyclic, thus enabling a reliable ranking to be extracted. By iteratively removing vertices based on their structural properties, we minimize the feedback arc set and ensure that the remaining graph is a DAG, which can be directly used to generate rankings in subsequent steps.
Algorithm 4 Preference Graphs Denoising for GED
1:$G$: Directed graph, var $s$: Vertex sequence
2:$s_{1}\leftarrow\emptyset$
3:$s_{2}\leftarrow\emptyset$
4:while $G\neq\emptyset$ do
5:     while $G$ contains a sink do
6:         Choose a sink $u$
7:         $s_{2}\leftarrow u\cdot s_{2}$
8:         $G\leftarrow G-u$
9:     end while
10:     while $G$ contains a source do
11:         Choose a source $u$
12:         $s_{1}\leftarrow s_{1}\cdot u$
13:         $G\leftarrow G-u$
14:     end while
15:     Choose a vertex $u$ with maximum $\delta(u)$
16:     $s_{1}\leftarrow s_{1}\cdot u$
17:     $G\leftarrow G-u$
18:end while
19:$s\leftarrow s_{1}\cdot s_{2}$
Appendix N Rank ensemble method
Weight score (Adler et al., 2002)
@@
Kemeny and weighted Kemeny (Kemeny, 1959)
@@
Pairwise majority and weighted pairwise majority (Caragiannis et al., 2016)
: The Pairwise Majority (PM) method determines a consensus ranking $\mathcal{R}^{*}$ by maximizing the number of pairwise agreements with the input rankings. For each pair of vertices $(v_{i},v_{j})$, the goal is to ensure that the majority of rankings agree with their relative order in $\mathcal{R}^{*}$:
$\mathcal{R}^{*}=\arg\max_{\mathcal{R}}\sum_{i<j}\left(\sum_{p=1}^{k}\mathds{1}(\mathcal{R}_{p}(v_{i})<\mathcal{R}_{p}(v_{j}))\right)\cdot\mathds{1}(\mathcal{R}(v_{i})<\mathcal{R}(v_{j}))$ (5)
@@
$\mathcal{R}^{*}=\arg\max_{\mathcal{R}}\sum_{i<j}\left(\sum_{p=1}^{k}\alpha_{p}\cdot\mathds{1}(\mathcal{R}_{p}(v_{i})<\mathcal{R}_{p}(v_{j}))\right)\cdot\mathds{1}(\mathcal{R}(v_{i})<\mathcal{R}(v_{j}))$ (6)
@@
Appendix O Visualization
(a) Raw Preference Graph 1
(b) Denoised Preference Graph 1
(c) Raw Preference Graph 2
(d) Denoised Preference Graph 2
Figure 5: Case studies showcasing the raw and denoised preference graphs. In both Case 1 and Case 2, the raw preference graphs (a, c) contain cyclic inconsistencies, which are resolved by GED into directed acyclic graphs (b, d). The dashed lines in the denoised graphs represent the edges that were removed by GED to eliminate cycles. The nodes labeled 0-9 correspond to the ten generated responses by Qwen2-72B. These examples illustrate the effectiveness of GED in eliminating noise and restoring consistency in preference evaluations.
In this section, we present two case studies that demonstrate the effectiveness of our proposed GED method in denoising preference graphs. Figure 5 illustrates the raw preference graphs, which are generated by multiple evaluators, i.e., Llama3-8B, Mistral-7B, and Qwen2-7B, through the responses produced by Qwen2-72B on the HumanEval benchmark. The nodes labeled 0-9 in the graphs correspond to the ten generated responses. The comparison between the raw (left) and denoised (right) graphs shows how our method successfully resolves cyclic inconsistencies, transforming noisy graphs into DAGs. In the denoised graphs, the dashed lines represent the edges that were removed by GED to eliminate cycles.
Case Study 1.
In the first case study, we showcase the impact of our denoising approach. The raw graph (Figure 5 (a)) contains multiple cyclic inconsistencies, such as $9\succ 4\succ 3\succ 9$, $9\succ 4\succ 3\succ 2\succ 9$, and $9\succ 4\succ 3\succ 7\succ 9$. By applying GED, we identify that removing a single edge (4 $\succ$ 3) eliminates all cycles, converting the noisy preference graph into a consistent DAG, as shown in Figure 5 (b).
Case Study 2.
The second case study (Figure 5 (c)) presents another scenario with conflicting preferences, where cycles like $7\succ 0\succ 6\succ 7$ and $7\succ 1\succ 0\succ 6\succ 7$ indicate noisy judgments. Here, removing just one edge (0 $\succ$ 6) using GED is sufficient to eliminate all cycles and convert the graph into a DAG, as depicted in Figure 5 (d).
Conclusion from Case Studies.
In both case studies, removing any other edge would not fully resolve all cyclic inconsistencies without requiring additional deletions, which would result in more information loss. GED effectively minimizes edge removals while maintaining the integrity of the original preference graph, making it a highly efficient solution for improving the consistency of preference evaluations.
Appendix P Prompt Template
@@
In this section, we provide detailed prompt templates used for response selection across five datasets: HumanEval (Chen et al., 2021), AlpacaEval (Dubois et al., 2024b), MATH (Hendrycks et al., 2021), GSM8k (Chen et al., 2021), and GAIA (Mialon et al., 2023). These prompts include both a system prompt to establish the evaluator’s context and a user prompt tailored to the specific task requirements. Each prompt is designed to guide the evaluators in comparing two candidate responses based on task-specific criteria such as correctness, clarity, efficiency, relevance, and completeness. The templates are shown in Table 11, Table 12, Table 13, Table 14, and Table 15.
@@
Prompt for Instruct Tuning.
In this section, we provide the prompt template used for data selection in the Instruct Tuning task. The goal is to select the most appropriate response for each instruction from multiple candidates, ensuring that the selected responses are helpful, harmless, and relevant. We use the HH-RLHF dataset (Bai et al., 2022), which contains human preference data aimed at training language models to be both helpful and harmless. The detailed prompt used by evaluators to assess and compare candidate responses is presented in Table 17.

### v4

@@ Appendix I The Importance of Addressing Cyclic Inconsistency
The selection of preference evaluators is a critical factor influencing the performance of PGED. While our initial experiments demonstrated the benefits of combining preference evaluators, further analysis is necessary to understand how their diversity and capabilities affect PGED’s effectiveness. To address this, we conducted additional experiments evaluating three different configurations of evaluator sets: the original set comprising Llama3-8B, Mistral-7B, and Qwen2-7B; a diverse set that adds Gemma-9B to introduce more variation in model architecture and training data; and a higher-capability set that replaces smaller models with larger ones, specifically Llama3-70B, Mistral-8×7B, and Qwen2-72B. The results, shown in Table 9, reveal two key insights. First, incorporating diversity by adding Gemma-9B improves PGED’s performance slightly, increasing the average score from 56.05% to 56.39%. This suggests that models with diverse training data and architectures can contribute to more robust aggregated evaluations. Second, replacing smaller evaluators with higher-capability models yields a more substantial improvement, with the average score rising to 60.89%. Notably, benchmarks like AlpacaEval and GAIA benefit the most from the advanced reasoning and language understanding capabilities of larger models. These findings demonstrate the importance of both diversity and capability in evaluator selection. Diversity provides marginal gains by bringing varied perspectives, while higher-capacity models contribute to more significant improvements by enhancing the overall quality of evaluations. This suggests that, when computational resources allow, incorporating advanced models into the evaluator set can meaningfully boost PGED’s performance.
@@
Table 9: Performance comparison with different evaluator sets across five benchmarks. The results highlight the impact of evaluator diversity and capability on PGED’s effectiveness.
@@
Appendix O Cost Considerations and Active Learning with ActiveGED
Using multiple evaluators and aggregating their preferences can be computationally expensive, especially when constructing dense preference graphs. This limitation becomes more pronounced in scenarios where preferences across all pairs of evaluators are required, which scales quadratically with the number of responses or models being compared. To address this, we clarify the computational trade-offs and propose an active learning-based approach to reduce the number of required pairwise evaluations. To further reduce the number of pairwise evaluations needed, we developed an active learning algorithm called ActiveGED. This algorithm strategically selects the most informative pairs for evaluation, effectively lowering the overall computational cost while maintaining high performance. ActiveGED combines random sampling with uncertainty-based selection to maximize information gain from each additional pairwise evaluation. We evaluated ActiveGED under budget constraints of 30% and 50% of the total possible pairwise comparisons, where we set the random budget ratio as 0.5. The results, presented in Table 10, demonstrate that ActiveGED achieves competitive performance with significantly fewer evaluations. For example, under a 50% budget, ActiveGED retains most of the performance benefits of full PGEDwhile cutting the number of pairwise comparisons in half.
@@
Evaluators Set of PGED HumanEval AlpacaEval MATH GSM8k GAIA Avg.
@@
ActiveGED demonstrates that by carefully selecting the most informative pairs, it is possible to achieve competitive performance with significantly fewer evaluations. This makes PGEDmore scalable and practical for real-world applications where computational resources are limited.
Appendix P Mitigating Evaluator Biases in PGED
Evaluator biases, such as position bias and token bias, can significantly impact the accuracy and fairness of the preference graphs used in PGED. Addressing these biases is crucial for ensuring reliable evaluations and robust performance. In this section, we describe the strategies employed in our framework to mitigate these biases and discuss potential areas for future improvement. Position bias arises when evaluators exhibit a preference for a particular position in a pairwise comparison, such as consistently favoring the first or second option regardless of content. To counter this, we explicitly include both orderings of each question and its candidate answers in the response selection setting. Specifically, for a question $Q$ with two candidate answers $A_{1}$ and $A_{2}$, we evaluate both configurations:
@@
By testing both (Q, A1, A2) and (Q, A2, A1), we ensure that any positional preferences of the evaluators are balanced out when constructing the preference graph. This approach minimizes the impact of position bias on the overall rankings. Token bias occurs when an evaluator exhibits a latent preference for a specific option or label (e.g., consistently favoring "Option A" over "Option B"). Our framework implicitly mitigates token bias through the aforementioned position-swapping strategy, which prevents evaluators from associating a fixed label with a specific position. By averaging the results across both orderings, any systematic preference for a particular label is effectively neutralized.
Appendix Q Rank ensemble method
Weight score (Adler, Friedman, and Sinuany-Stern 2002)
@@
Kemeny and weighted Kemeny (Kemeny 1959)
@@
Pairwise majority and weighted pairwise majority
(Caragiannis, Procaccia, and Shah 2016): The Pairwise Majority (PM) method determines a consensus ranking $\mathcal{R}^{*}$ by maximizing the number of pairwise agreements with the input rankings. For each pair of vertices $(v_{i},v_{j})$, the goal is to ensure that the majority of rankings agree with their relative order in $\mathcal{R}^{*}$:
$\mathcal{R}^{*}=\arg\max_{\mathcal{R}}\sum_{i<j}\left(\sum_{p=1}^{k}\mathds{1}(\mathcal{R}_{p}(v_{i})<\mathcal{R}_{p}(v_{j}))\right)\cdot\mathds{1}(\mathcal{R}(v_{i})<\mathcal{R}(v_{j}))$
@@
$\mathcal{R}^{*}=\arg\max_{\mathcal{R}}\sum_{i<j}\left(\sum_{p=1}^{k}\alpha_{p}\cdot\mathds{1}(\mathcal{R}_{p}(v_{i})<\mathcal{R}_{p}(v_{j}))\right)\cdot\mathds{1}(\mathcal{R}(v_{i})<\mathcal{R}(v_{j}))$
@@
Appendix R Cost-Performance Analysis of PGED
We conducted a detailed cost-performance analysis based on publicly available API pricing from Amazon Bedrock. API Cost Comparison: For our PGED method, we utilize three relatively lightweight evaluators: Llama3-8B, Mistral-7B, and Qwen2-7B. Since pricing information for Qwen2 is not currently available on Amazon Bedrock, we conservatively adopt the pricing of Llama3 as a proxy. Specifically, the cost per 1 million input tokens is \$0.09, and the cost per 1 million output tokens is \$0.175. In contrast, using a single strong evaluator such as Qwen2-72B is significantly more expensive, with costs of \$0.23 per 1 million input tokens and \$0.40 per 1 million output tokens. This comparison shows that our PGED method can achieve superior evaluation performance at only approximately 39.13% of the input cost and 43.75% of the output cost. Performance vs. Budget Trade-off: As demonstrated in Appendix O, PGED is capable of outperforming strong evaluators like Qwen2-72B even when utilizing only 30–50% of the evaluation budget. This efficiency arises from effectively aggregating and denoising multiple weaker signals. The computational overhead involved in graph fusion and denoising steps within PGED is minimal compared to the inference time of large language models, making PGED a cost-effective and practical solution for scalable preference evaluations.
Table 11: Performance comparison of PGEDwith multi-evaluator baselines PRD and ChatEval. PGEDachieves the highest score on each dataset and the best overall average.
Method HumanEval AlpacaEval MATH GSM8k GAIA Avg.
PRD 66.53 30.33 74.89 86.75 11.83 54.07
ChatEval 66.42 29.71 74.62 86.66 11.53 53.79
PGED 70.73 32.44 75.58 88.18 13.33 56.05
Appendix S Comparison with More Baselines
We extend our experimental analysis to include two recent multi-evaluator baselines: PRD (Li, Patel, and Du 2023b) and ChatEval (Chan et al. 2023b). Both methods are designed to improve evaluation reliability by aggregating feedback from multiple evaluators. PRD focuses on pairwise ranking distillation, while ChatEval employs evaluator voting mechanisms. For a fair comparison, we evaluate all methods using the same set of models: Llama3-8B, Mistral-7B, and Qwen2-7B. Table 11 summarizes the results across five standard benchmark datasets. PGEDconsistently outperforms both PRD and ChatEval across all tasks. These results highlight the effectiveness of PGED’s structural approach. Unlike debate-style or voting-based methods, PGEDconstructs and denoises preference graphs to enforce global consistency, resulting in both improved empirical performance and stronger theoretical guarantees.
Appendix T PGED with Strong Evaluators
To examine whether PGED can benefit from stronger evaluators, we conducted additional experiments using three high-capacity language models: GPT-4-o-mini, GPT-4-o, and Gemini 2.0. These evaluators were used to construct individual and ensemble preference graphs under the PGED framework. The results, presented in Figure 6, demonstrate clear improvements of PGED over each individual evaluator across multiple tasks. This confirms that PGED not only performs well with lightweight models but also enhances the evaluation quality when applied to strong LLMs such as GPT-4 and Gemini. These findings support the generality and scalability of our framework. By aggregating and denoising even high-quality but noisy judgments, PGED consistently improves robustness and overall evaluation accuracy.
Figure 6: Performance of PGED using strong evaluators: GPT-4-o-mini, GPT-4-o, and Gemini 2.0. PGED ensemble improves upon each individual model across tasks.
Appendix U Comparison with Weighted Voting Baselines
We further evaluate PGED against two intuitive multi-evaluator aggregation baselines: Weighted Majority Voting and Weighted Average of Ratings. Both baselines utilize the same set of evaluators (Llama3-8B, Mistral-7B, and Qwen2-7B) and apply weight-based schemes to either majority votes or numerical scores. Weighted Majority Voting assigns each evaluator a weight and determines the preferred candidate based on weighted voting outcomes across all pairwise comparisons. Weighted Average of Ratings aggregates normalized point-wise scores by evaluator weights to rank responses. As shown in Figure 7, PGED consistently outperforms both baselines across all evaluated tasks. This demonstrates that leveraging the structure of preference graphs and enforcing global consistency through denoising leads to more accurate and robust evaluations compared to simple weighted heuristics.
Figure 7: Performance comparison of GED with Weighted Majority Voting and Weighted Average of Ratings. GED outperforms both across tasks.
Appendix V Prompt Template
@@
In this section, we provide detailed prompt templates used for response selection across five datasets: HumanEval (Chen et al. 2021), AlpacaEval (Dubois et al. 2024b), MATH (Hendrycks et al. 2021), GSM8k (Chen et al. 2021), and GAIA (Mialon et al. 2023). These prompts include both a system prompt to establish the evaluator’s context and a user prompt tailored to the specific task requirements. Each prompt is designed to guide the evaluators in comparing two candidate responses based on task-specific criteria such as correctness, clarity, efficiency, relevance, and completeness. The templates are shown in Table 12, Table 13, Table 14, Table 15, and Table 16.
Prompt for Model Ranking.
Prompt for Instruct Tuning.
In this section, we provide the prompt template used for data selection in the Instruct Tuning task. The goal is to select the most appropriate response for each instruction from multiple candidates, ensuring that the selected responses are helpful, harmless, and relevant. We use the HH-RLHF dataset (Bai et al. 2022), which contains human preference data aimed at training language models to be both helpful and harmless. The detailed prompt used by evaluators to assess and compare candidate responses is presented in Table 18.